• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name kelleher

    Log entry time 11:05:58 on July11,2010

    Entry number 331450

    keyword=back to 120uA

    For a brief period starting at about 09:40, MCC sent 100uA of beam. This was due to trips in the beam loss monitor (blm). This appears to have been corrected as we are back to running at 120uA.


    FIGURE 1